# Training-Free Bias Mitigation by LLM-Assisted Bias Detection and Latent Variable Guidance

### v3

@@
Training-Free Bias Mitigation by LLM-Assisted Bias Detection and Latent Variable Guidance
Jinya Sakurai,  Yuki Koyama,  Issei Sato
The University of Tokyo
sakurai-jinya725@g.ecc.u-tokyo.ac.jp,   koyama@pe.t.u-tokyo.ac.jp,
sato@g.ecc.u-tokyo.ac.jp
Corresponding author: sakurai-jinya725@g.ecc.u-tokyo.ac.jp
@@ 1 Introduction
Text-to-image (T2I) models, such as Stable Diffusion (sd3; sdxl), Imagen (imagen3), and Flux (flux), have transformed visual content creation by allowing users to generate high-quality images from natural language descriptions. Yet, despite their accessibility and creative potential, these models inherit and amplify the biases present in large-scale web datasets (bavalatti2025systematic; naik2023social). When prompted with neutral descriptions, models often produce stereotypical representations such as depicting a nurse as female or a chief executive officer as male, thereby reinforcing societal inequalities (wan2024survey; mandal2024generated; sandoval2024perpetuation; zhang2024partiality; dealmeida2024bias). The increasing reuse of synthetic images in training pipelines further intensifies these biases over time (wyllie2024fairness; alemohammad2023self).
Many algorithmic approaches have been proposed to mitigate this problem; however, they continue to face significant practical limitations from the user’s perspective. Large-scale dataset curation is challenging to sustain over time (bianchi2023easily), and most existing strategies (Kim_2024; zhang2023iti; hirota2024saner) rely on computationally expensive fine-tuning or retraining for each target bias, often involving delicate hyperparameter tuning (friedrich2023fair). In addition, these approaches typically depend on manually pre-defined attribute lists (Kim_2024; zhang2023iti; friedrich2023fair), which not only constrains their adaptability to emerging or context-dependent biases but also limits the degree of user control and interpretability during deployment.
@@ 2 Related Work
The generative capabilities of T2I models (imagen3; sd3; chen2025pixart; flux) have improved dramatically through several key developments: training on large scale text image pairs (schuhmann2022laion; Chen_2023), architectural improvements (peebles2023scalable) from UNet (ronneberger2015u) to transformer-based designs (AttentionAllYouNeed; Dosovitskiy_2020), and theoretical developments from diffusion models (sohl2015deep; ddpm; Song_2020) to flow matching (Lipman_2022; Xingchao_2022; lipman2024flowmatchingguidecode). Furthermore, recent fine tuning approaches have significantly reduced generation time by minimizing the required number of inference steps (liu2023instaflow; chen2024pixartdeltafastcontrollableimage; sauer2024fast; Sauer_2023; Yang_2023). A distinctive characteristic of state- of-the-art models is their use of separately trained components, such as Variational Auto Encoders (kingma_2014) for image compression, text encoders (raffel2020exploring; clip), and flow model backbones, each trained on different datasets.
@@
Numerous studies have investigated biases in T2I models. Research such as Ghosh_2023; Wu_2023; bianchi2023easily highlights the biases embedded in generated outputs for seemingly neutral input prompts that lack explicit identity- or demographic-related terms. Other works, including Cho_2022; Luccioni_2023 predefine sensitive human attributes and analyze biases in outputs generated from occupational input prompts. naik2023social provides broader analyses, including comparative studies with statistical data or image search results, as well as spatial analyses of generated images. Wang_2023 applies methods from social psychology to explore implicit and complex biases related to race and gender. Furthermore, Luccioni_2023 introduces an interactive bias analysis tool leveraging clustering methods. Lastly, Chen_2024 examines the potential for AI-generated images to perpetuate harmful feedback loops, amplifying biases in AI systems when used as training data for future models. D'Incà_2024; Chinchure_2023 employ LLMs to detect open-ended biases in text-to-image models where users do not have to provide predefined bias attributes.
@@
One line of work focuses on making T2I models fair by fine-tuning or retraining various components, from the diffusion backbone (Kim_2024) and text encoder hirota2024saner to postfix prompt embeddings zhang2023iti or additive residual image vectors Seth_2023_CVPR. While effective, such retraining is both time- and resource-intensive and must be repeated whenever societal notions of bias evolve. By contrast, several methods eliminate societal bias only by steering generation at inference time. FairDiffusion friedrich2023fair guides sampling with randomly chosen bias attributes; (chuang2023debiasing) removes unwanted directions from the text embedding space via calibrated projection matrices; and ENTIGEN bansal2022well simply appends an ethical fairness postfix to the input prompt. Although these approaches require no additional training, they do depend on a predefined list of bias categories and attribute values for control.
@@
Recent research has increasingly explored user interfaces that support interactive control and understanding of T2I generation. PromptCharm (wang2024promptcharm) introduced a mixed initiative interface that refines prompts through multi-modal feedback and visual explanations, allowing users to iteratively steer image generation with system suggestions. Promptify (brade2023promptify) similarly employed LLMs to recommend prompt modifications based on user feedback, enabling interactive exploration and refinement of text prompts. Beyond linguistic prompting, PromptPaint (chung2023promptpaint) introduced an interactive canvas that enables users to spatially compose scenes by assigning distinct text prompts to different regions of the image. PromptMap (adamkiewicz2025promptmap) shifted the focus from control to sensemaking, visualizing the branching structure of prompt evolution as a map to help users navigate and reflect on their creative process. Expandora (choi2025expandora) extended these ideas to structured exploration, allowing users to specify design intentions and desired diversity levels to automatically generate a wide range of prompt variations. Collectively, these systems emphasize human agency, transparency, and creativity in T2I interaction. In contrast, FairT2I focuses on societal bias detection and mitigation, offering an interface that reveals and allows direct manipulation of latent bias attributes, thereby bridging algorithmic fairness and interactive generative control.
@@ 3.1 Preliminaries: Classifier-Free Guidance
In diffusion models (ddpm; sohl2015deep) and flow matching models (sd3; Lipman_2022), classifier-free guidance (ho2022classifier) is a powerful tool for sampling images $\mathbf{x}$ conditioned on a text input $\mathbf{y}$. It accomplishes this by guiding an unconditional image generation model using the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. This enables the sampling from
@@ 3.2 Latent Variable Guidance for Bias Mitigation
This formulation, presented in Equation (1), effectively decomposes the score function guiding the image generation process. It explicitly separates the influence of the latent attribute $\mathbf{z}$ (which is sampled according to its conditional probability $p(\mathbf{z}\mid\mathbf{y})$) from the subsequent image generation step, which is guided by both this sampled attribute $\mathbf{z}$ and the original input text $\mathbf{y}$. This perspective allows us to contextualize existing bias mitigation strategies. Approaches such as replacing text encoders with fair alternatives (hirota2024saner; chuang2023debiasing), learning fair prompt embeddings (zhang2023iti), or randomly sampling sensitive attributes (friedrich2023fair) can all be interpreted as methods that implicitly or explicitly substitute the T2I model’s original, potentially biased distribution $p_{\text{bias}}(\mathbf{z}\mid\mathbf{y})$ with a more equitable target distribution $p_{\text{fair}}(\mathbf{z}\mid\mathbf{y})$. Figure 2 illustrates the generation process with latent variable guidance, compared to the standard process without it, for an input prompt $\mathbf{y}=\textit{A CEO}$ and an attribute set $\mathcal{Z}=\{z_{1}=\textit{male},z_{2}=\textit{female}\}$. In the standard setting, images are denoised at each timestep under the biased distribution $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$, which tends to steer the generation toward male CEO representations. In contrast, latent variable guidance enables the replacement of $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$ with a fair distribution $p_{\mathrm{fair}}(\mathbf{z}\mid\mathbf{y})$, allowing the generation process to be steered toward more balanced and fair outcomes.
@@ 3.3 LLM-Assisted Bias Detection
To address these challenges, we leverage large language models (LLMs) (gpt4; anthropic2025claude37sonnet) to automatically detect biases in the input text, as with existing bias detection methods (D'Incà_2024; Chinchure_2023). Specifically, we use the LLM to predict the set of possible latent attributes $\mathcal{Z}$ from the input text $\mathbf{y}$. LLMs are prompted to output a set of latent attributes that are likely to appear in images generated by T2I models with the input text in a JSON format. This approach allows us to handle a broader range of input prompts and to detect biases that may not be apparent to human annotators.
@@ 3.4 Attribute Resampling
One simple approach is to set the distribution of latent attributes to be uniform across all possible values of $z$: $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})=\frac{1}{|\mathcal{Z}|}$, which corresponds to simply mixing the scores $\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$ with equal proportions across all possible values of $z$. This formulation coincides with that of Fair Diffusion (friedrich2023fair).
@@
Research has shown that T2I models tend to exaggerate demographic stereotypes beyond what we observe in real-world distributions across various latent attributes (naik2023social). One way to address this issue is to incorporate real-world statistical data as $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$, ensuring that the generated image distributions are at least as balanced as real-world demographics. [^1] It is worth noting that real-world occupational distributions often reflect systemic biases and unequal access to opportunities, shaped by historical and societal factors such as limited access to education or workplace discrimination. These disparities highlight that real-world distributions are not inherently fair.
@@ 4.1 Experimental Settings
We used two datasets for this study: one for assessing societal biases in professions and another for examining potential biases in more diverse textual prompts. To evaluate occupational bias, we used the Stable Bias profession dataset (Luccioni_2023), which contains $131$ occupations collected from the U.S. Bureau of Labor Statistics (BLS) (see Appendix A of Luccioni_2023 for full details). To probe broader forms of bias, we adopt the Parti Prompt dataset (yu2022scaling), which includes more than $1,600$ English prompts designed to evaluate text-to-image generation models and test their limitations. For efficiency, we randomly sample $50$ prompts from each dataset, with the selected subset listed in subsection B.1.
@@
We used Claude 3.7-sonnet for LLM-assisted bias detection, following a small pilot study that compared GPT-4o gpt4, Claude 3.7-sonnet anthropic2025claude37sonnet, Llama 3.3 grattafiori2024llama3herdmodels, and DeepSeek V3 liu2024deepseek. Further details of this pilot study are provided in the appendix Appendix B.
@@ 4.2 Human Annotation and Reference Set Construction
To establish a human reference for bias categories, we conducted an annotation study using the same prompts as in the evaluation datasets. 100 text prompts, consisting of 50 prompts from the Stable Bias dataset (Luccioni_2023) and $50$ prompts from the PartiPrompts dataset (yu2022scaling) were divided into five groups of $20$ prompts, with each group assigned to five independent annotators ($25$ participants in total). Annotators were instructed to list potential bias categories (e.g., gender, race, attire) along with multiple attributes per category (e.g., male, female, non-binary for gender). Clear format guidelines and illustrative examples were provided to ensure consistency. To maintain parity, the same few-shot examples were used both in the human instructions and in the system prompts provided to the LLMs.
@@ 5.1 Automatic Evaluation
To evaluate societal bias in professions, we use the same Stable Bias profession dataset (Luccioni_2023) introduced in section 4, which contains 131 occupations from the U.S. Bureau of Labor Statistics (BLS). We consider two target distributions: (1) the uniform distribution over categories defined in FairFace (karkkainen2021fairface), and (2) BLS occupational statistics. For the uniform distribution, we used the same subset of 50 randomly sampled occupations from this dataset as in section 4. The selected subset of prompts is listed in subsection B.1. For the BLS statistics, we use the five occupation prompts mentioned in naik2023social: CEO, doctor, computer programmer, nurse, and house keeper. As in zhang2023iti, each profession is prefixed with “A headshot of ” to form an input prompt.
@@
Following hirota2024saner; chuang2023debiasing, we measure bias in the generated outputs by computing the statistical parity (choi2020fair) between the empirical distribution and the ideal target distribution. In line with naik2023social, we detect faces in the generated images using dlib (king2015max) and then predict the attributes of each image using classifiers trained on the FairFace dataset (karkkainen2021fairface). Statistical parity is defined as the $\ell_{2}$‑distance between two probability vectors. Specifically, let $\mathbf{p}=(p_{1},p_{2},\dots,p_{n}),\quad\mathbf{q}=(q_{1},q_{2},\dots,q_{n})$ be two distributions over $n$ outcomes. Then the statistical parity between $\mathbf{p}$ and $\mathbf{q}$ is
@@
  • Original: the unmodified Stable Diffusion model (sd), with no bias mitigation.
  • ENTIGEN (bansal2022well): appending an ethical injection phrase to the end of the prompt.
  • Fair Diffusion (friedrich2023fair): randomly sampling attributes from a predefined bias set and applying semantic guidance (brack2023sega).
Detailed hyperparameter settings for each method are provided in subsection C.2 and subsection C.3, respectively. To ensure a fair comparison, the T2I model was standardized to Stable Diffusion 1.5 (rombach2022high) across all baseline and proposed methods.
@@
We used Stable Diffusion 1.5 (rombach2022high) as our T2I model. This choice was made because the hyperparameter settings and text encoders of the comparison methods were not compatible with the newer T2I architectures (sd3). We employed Claude 3.7-sonnet (anthropic2025claude37sonnet) for LLM-assisted bias detection, and GPT4o-mini (gpt4) to fuse the textual input $\bf{y}$ with the sensitive attribute $\bf{z}$. The parameter settings for image generation and the exact prompts supplied to the LLM are detailed in subsection C.4.
@@
ENTIGEN (bansal2022well) 0.3748 0.0569 0.1229 0.1210 0.0121
FairDiffusion (friedrich2023fair) 0.1956 0.1089 0.1138 0.1653 0.1597
@@
ENTIGEN (bansal2022well) 0.5004 0.6115 0.4727 0.4805 0.9341
FairDiffusion (friedrich2023fair) 0.1664 0.1484 0.0750 0.1285 0.3074

### v4

@@
Abstract
Text-to-image (T2I) models have advanced creative content generation, yet their reliance on large uncurated datasets often reproduces societal biases. We present FairT2I, a training-free and interactive framework grounded in a mathematically principled latent variable guidance formulation. This formulation decomposes the generative score function into attribute-conditioned components and reweights them according to a defined distribution, providing a unified and flexible mechanism for bias-aware generation that also subsumes many existing ad hoc debiasing approaches as special cases. Building upon this foundation, FairT2I incorporates (1) latent variable guidance as the core mechanism, (2) LLM-based bias detection to automatically infer bias-prone categories and attributes from text prompts as part of the latent structure, and (3) attribute resampling, which allows users to adjust or redefine the attribute distribution based on uniform, real-world, or user-specified statistics. The accompanying user interface supports this pipeline by enabling users to inspect detected biases, modify attributes or weights, and generate debiased images in real time. Experimental results show that LLMs outperform average human annotators in the number and granularity of detected bias categories and attributes. Moreover, FairT2I achieves superior performance to baseline models in both societal bias mitigation and image diversity, while preserving image quality and prompt fidelity.
@@ 1 Introduction
Text-to-image (T2I) models, such as Stable Diffusion (Esser et al., 2024; Podell et al., 2023), Imagen (Imagen-Team-Google et al., 2024), and Flux (Black Forest Labs, 2024), have transformed visual content creation by allowing users to generate high-quality images from natural language descriptions. Yet, despite their accessibility and creative potential, these models inherit and amplify the biases present in large-scale web datasets (Bavalatti et al., 2025; Naik and Nushi, 2023). When prompted with neutral descriptions, models often produce stereotypical representations such as depicting a nurse as female or a chief executive officer as male, thereby reinforcing societal inequalities (Wan et al., 2024; Mandal et al., 2024; Sandoval-Martin and Martínez-Sanzo, 2024; Zhang et al., 2024b; de Almeida and Rafael, 2024). The increasing reuse of synthetic images in training pipelines further intensifies these biases over time (Wyllie et al., 2024; Alemohammad et al., 2023).
Many algorithmic approaches have been proposed to mitigate this problem; however, they continue to face significant practical limitations from the user’s perspective. Large-scale dataset curation is challenging to sustain over time (Bianchi et al., 2023), and most existing strategies (Kim et al., 2024; Zhang et al., 2023; Hirota et al., 2024) rely on computationally expensive fine-tuning or retraining for each target bias, often involving delicate hyperparameter tuning (Friedrich et al., 2023). In addition, these approaches typically depend on manually pre-defined attribute lists (Kim et al., 2024; Zhang et al., 2023; Friedrich et al., 2023), which not only constrains their adaptability to emerging or context-dependent biases but also limits the degree of user control and interpretability during deployment.
@@ 2 Related Work
The generative capabilities of T2I models (Imagen-Team-Google et al., 2024; Esser et al., 2024; Chen et al., 2025a; Black Forest Labs, 2024) have improved dramatically through several key developments: training on large scale text image pairs (Schuhmann et al., 2022; Chen et al., 2023), architectural improvements (Peebles and Xie, 2023) from UNet (Ronneberger et al., 2015) to transformer-based designs (Vaswani et al., 2017; Dosovitskiy et al., 2020), and theoretical developments from diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) to flow matching (Lipman et al., 2022; Liu et al., 2022; Lipman et al., 2024). Furthermore, recent fine tuning approaches have significantly reduced generation time by minimizing the required number of inference steps (Liu et al., 2023; Chen et al., 2024a; Sauer et al., 2024, 2023; Yang et al., 2023). A distinctive characteristic of state- of-the-art models is their use of separately trained components, such as Variational Auto Encoders (Kingma et al., 2014) for image compression, text encoders (Raffel et al., 2020; Radford et al., 2021), and flow model backbones, each trained on different datasets.
@@
Numerous studies have investigated biases in T2I models. Research such as Ghosh and Caliskan (2023); Wu et al. (2023); Bianchi et al. (2023) highlights the biases embedded in generated outputs for seemingly neutral input prompts that lack explicit identity- or demographic-related terms. Other works, including Cho et al. (2022); Luccioni et al. (2023) predefine sensitive human attributes and analyze biases in outputs generated from occupational input prompts. Naik and Nushi (2023) provides broader analyses, including comparative studies with statistical data or image search results, as well as spatial analyses of generated images. Wang et al. (2023) applies methods from social psychology to explore implicit and complex biases related to race and gender. Furthermore, Luccioni et al. (2023) introduces an interactive bias analysis tool leveraging clustering methods. Lastly, Chen et al. (2024b) examines the potential for AI-generated images to perpetuate harmful feedback loops, amplifying biases in AI systems when used as training data for future models. D'Incà_2024; Chinchure et al. (2023) employ LLMs to detect open-ended biases in text-to-image models where users do not have to provide predefined bias attributes.
@@
One line of work focuses on making T2I models fair by fine-tuning or retraining various components, from the diffusion backbone (Kim et al., 2024) and text encoder Hirota et al. (2024) to postfix prompt embeddings Zhang et al. (2023) or additive residual image vectors Seth et al. (2023). While effective, such retraining is both time- and resource-intensive and must be repeated whenever societal notions of bias evolve. By contrast, several methods eliminate societal bias only by steering generation at inference time. FairDiffusion Friedrich et al. (2023) guides sampling with randomly chosen bias attributes; (Chuang et al., 2023) removes unwanted directions from the text embedding space via calibrated projection matrices; and ENTIGEN Bansal et al. (2022) simply appends an ethical fairness postfix to the input prompt. Although these approaches require no additional training, they do depend on a predefined list of bias categories and attribute values for control.
@@
Recent research has increasingly explored user interfaces that support interactive control and understanding of T2I generation. PromptCharm (Wang et al., 2024) introduced a mixed initiative interface that refines prompts through multi-modal feedback and visual explanations, allowing users to iteratively steer image generation with system suggestions. Promptify (Brade et al., 2023) similarly employed LLMs to recommend prompt modifications based on user feedback, enabling interactive exploration and refinement of text prompts. Beyond linguistic prompting, PromptPaint (Chung and Adar, 2023) introduced an interactive canvas that enables users to spatially compose scenes by assigning distinct text prompts to different regions of the image. PromptMap (Adamkiewicz et al., 2025) shifted the focus from control to sensemaking, visualizing the branching structure of prompt evolution as a map to help users navigate and reflect on their creative process. Expandora (Choi et al., 2025) extended these ideas to structured exploration, allowing users to specify design intentions and desired diversity levels to automatically generate a wide range of prompt variations. Collectively, these systems emphasize human agency, transparency, and creativity in T2I interaction. In contrast, FairT2I focuses on societal bias detection and mitigation, offering an interface that reveals and allows direct manipulation of latent bias attributes, thereby bridging algorithmic fairness and interactive generative control.
@@ 3.1 Preliminaries: Classifier-Free Guidance
In diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) and flow matching models (Esser et al., 2024; Lipman et al., 2022), classifier-free guidance (Ho and Salimans, 2022) is a powerful tool for sampling images $\mathbf{x}$ conditioned on a text input $\mathbf{y}$. It accomplishes this by guiding an unconditional image generation model using the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. This enables the sampling from
@@ 3.2 Latent Variable Guidance for Bias Mitigation
This formulation, presented in Equation (1), effectively decomposes the score function guiding the image generation process. It explicitly separates the influence of the latent attribute $\mathbf{z}$ (which is sampled according to its conditional probability $p(\mathbf{z}\mid\mathbf{y})$) from the subsequent image generation step, which is guided by both this sampled attribute $\mathbf{z}$ and the original input text $\mathbf{y}$. This perspective allows us to contextualize existing bias mitigation strategies. Approaches such as replacing text encoders with fair alternatives (Hirota et al., 2024; Chuang et al., 2023), learning fair prompt embeddings (Zhang et al., 2023), or randomly sampling sensitive attributes (Friedrich et al., 2023) can all be interpreted as methods that implicitly or explicitly substitute the T2I model’s original, potentially biased distribution $p_{\text{bias}}(\mathbf{z}\mid\mathbf{y})$ with a more equitable target distribution $p_{\text{fair}}(\mathbf{z}\mid\mathbf{y})$. Figure 2 illustrates the generation process with latent variable guidance, compared to the standard process without it, for an input prompt $\mathbf{y}=\textit{A CEO}$ and an attribute set $\mathcal{Z}=\{z_{1}=\textit{male},z_{2}=\textit{female}\}$. In the standard setting, images are denoised at each timestep under the biased distribution $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$, which tends to steer the generation toward male CEO representations. In contrast, latent variable guidance enables the replacement of $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$ with a fair distribution $p_{\mathrm{fair}}(\mathbf{z}\mid\mathbf{y})$, allowing the generation process to be steered toward more balanced and fair outcomes.
@@ 3.3 LLM-Assisted Bias Detection
To address these challenges, we leverage large language models (LLMs) (OpenAI et al., 2024; Anthropic, 2025) to automatically detect biases in the input text, as with existing bias detection methods (D'Incà_2024; Chinchure et al., 2023). Specifically, we use the LLM to predict the set of possible latent attributes $\mathcal{Z}$ from the input text $\mathbf{y}$. LLMs are prompted to output a set of latent attributes that are likely to appear in images generated by T2I models with the input text in a JSON format. This approach allows us to handle a broader range of input prompts and to detect biases that may not be apparent to human annotators.
@@ 3.4 Attribute Resampling
One simple approach is to set the distribution of latent attributes to be uniform across all possible values of $z$: $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})=\frac{1}{|\mathcal{Z}|}$, which corresponds to simply mixing the scores $\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$ with equal proportions across all possible values of $z$. This formulation coincides with that of Fair Diffusion (Friedrich et al., 2023).
@@
Research has shown that T2I models tend to exaggerate demographic stereotypes beyond what we observe in real-world distributions across various latent attributes (Naik and Nushi, 2023). One way to address this issue is to incorporate real-world statistical data as $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$, ensuring that the generated image distributions are at least as balanced as real-world demographics. [^1] It is worth noting that real-world occupational distributions often reflect systemic biases and unequal access to opportunities, shaped by historical and societal factors such as limited access to education or workplace discrimination. These disparities highlight that real-world distributions are not inherently fair.
@@ 4.1 Experimental Settings
We used two datasets for this study: one for assessing societal biases in professions and another for examining potential biases in more diverse textual prompts. To evaluate occupational bias, we used the Stable Bias profession dataset (Luccioni et al., 2023), which contains $131$ occupations collected from the U.S. Bureau of Labor Statistics (BLS) (see Appendix A of Luccioni et al. (2023) for full details). To probe broader forms of bias, we adopt the Parti Prompt dataset (Yu et al., 2022), which includes more than $1,600$ English prompts designed to evaluate text-to-image generation models and test their limitations. For efficiency, we randomly sample $50$ prompts from each dataset, with the selected subset listed in subsection B.1.
@@
We used Claude 3.7-sonnet for LLM-assisted bias detection, following a small pilot study that compared GPT-4o OpenAI et al. (2024), Claude 3.7-sonnet Anthropic (2025), Llama 3.3 Grattafiori et al. (2024), and DeepSeek V3 Liu et al. (2024). Further details of this pilot study are provided in the appendix Appendix B.
@@ 4.2 Human Annotation and Reference Set Construction
To establish a human reference for bias categories, we conducted an annotation study using the same prompts as in the evaluation datasets. 100 text prompts, consisting of 50 prompts from the Stable Bias dataset (Luccioni et al., 2023) and $50$ prompts from the PartiPrompts dataset (Yu et al., 2022) were divided into five groups of $20$ prompts, with each group assigned to five independent annotators ($25$ participants in total). Annotators were instructed to list potential bias categories (e.g., gender, race, attire) along with multiple attributes per category (e.g., male, female, non-binary for gender). Clear format guidelines and illustrative examples were provided to ensure consistency. To maintain parity, the same few-shot examples were used both in the human instructions and in the system prompts provided to the LLMs.
@@ 5.1 Automatic Evaluation
To evaluate societal bias in professions, we use the same Stable Bias profession dataset (Luccioni et al., 2023) introduced in section 4, which contains 131 occupations from the U.S. Bureau of Labor Statistics (BLS). We consider two target distributions: (1) the uniform distribution over categories defined in FairFace (Karkkainen and Joo, 2021), and (2) BLS occupational statistics. For the uniform distribution, we used the same subset of 50 randomly sampled occupations from this dataset as in section 4. The selected subset of prompts is listed in subsection B.1. For the BLS statistics, we use the five occupation prompts mentioned in Naik and Nushi (2023): CEO, doctor, computer programmer, nurse, and house keeper. As in Zhang et al. (2023), each profession is prefixed with “A headshot of ” to form an input prompt.
@@
Following Hirota et al. (2024); Chuang et al. (2023), we measure bias in the generated outputs by computing the statistical parity (Choi et al., 2020) between the empirical distribution and the ideal target distribution. In line with Naik and Nushi (2023), we detect faces in the generated images using dlib (King, 2015) and then predict the attributes of each image using classifiers trained on the FairFace dataset (Karkkainen and Joo, 2021). Statistical parity is defined as the $\ell_{2}$‑distance between two probability vectors. Specifically, let $\mathbf{p}=(p_{1},p_{2},\dots,p_{n}),\quad\mathbf{q}=(q_{1},q_{2},\dots,q_{n})$ be two distributions over $n$ outcomes. Then the statistical parity between $\mathbf{p}$ and $\mathbf{q}$ is
@@
  • Original: the unmodified Stable Diffusion model (Rombach et al., 2022a), with no bias mitigation.
  • ENTIGEN (Bansal et al., 2022): appending an ethical injection phrase to the end of the prompt.
  • Fair Diffusion (Friedrich et al., 2023): randomly sampling attributes from a predefined bias set and applying semantic guidance (Brack et al., 2023).
Detailed hyperparameter settings for each method are provided in subsection C.2 and subsection C.3, respectively. To ensure a fair comparison, the T2I model was standardized to Stable Diffusion 1.5 (Rombach et al., 2022b) across all baseline and proposed methods.
@@
We used Stable Diffusion 1.5 (Rombach et al., 2022b) as our T2I model. This choice was made because the hyperparameter settings and text encoders of the comparison methods were not compatible with the newer T2I architectures (Esser et al., 2024). We employed Claude 3.7-sonnet (Anthropic, 2025) for LLM-assisted bias detection, and GPT4o-mini (OpenAI et al., 2024) to fuse the textual input $\bf{y}$ with the sensitive attribute $\bf{z}$. The parameter settings for image generation and the exact prompts supplied to the LLM are detailed in subsection C.4.
@@
ENTIGEN (Bansal et al., 2022) 0.3748 0.0569 0.1229 0.1210 0.0121
FairDiffusion (Friedrich et al., 2023) 0.1956 0.1089 0.1138 0.1653 0.1597
@@
ENTIGEN (Bansal et al., 2022) 0.5004 0.6115 0.4727 0.4805 0.9341
FairDiffusion (Friedrich et al., 2023) 0.1664 0.1484 0.0750 0.1285 0.3074